OpenStudy (anonymous):

3(2x – 7) = 3 What are the solutions to this equation? Show your work.

OpenStudy (yoloshroom):

multiply it out, so 3*2x +3*-7 = 3 then move x to one side and divide so X is on it's own

OpenStudy (anonymous):

3(2x-7) = 3 6x-21=3 6x-24=0 x=4

OpenStudy (yoloshroom):

Comon jayyy we want them to do theiir owwwn work :(

OpenStudy (anonymous):

why multiply out?

OpenStudy (anonymous):

its fine im ok with anything :) I just need to finish soon

OpenStudy (anonymous):

divide by 3 and start with \[2x-7=1\]